Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,100 units × $20 per unit) $ 262,000
Variable expenses 131,000
Contribution margin 131,000
Fixed expenses 146,000
Net operating loss $ (15,000 )

Required:

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $56,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,100 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,100)?
Answer #1
5
a. Current variable expense per unit=Total variable expenses/Units sold=131000/13100=$ 10 per unit
Revised variable expenses per unit=10-3=$ 7 per unit
Revised fixed expenses=146000+56000=$ 202000
CM ratio=Contribution margin/Selling price per unit
Contribution margin=Selling price-Variable expenses=20-7=$ 13
CM ratio=13/20=0.65
Break-even point in unit sales=Fixed expenses/Contribution margin per unit=202000/13=15538.46=15539 units
Break-even point in dollar sales=Fixed expenses/Contribution margin ratio=202000/0.65=$ 310769
b. Contribution format income statement:
Not automated Automated
Per unit % Total Per unit % Total
Sales 20 100% 402000 20 100% 402000
(20100*20) (20100*20)
Less: Variable expenses 10 50% 201000 7 35% 140700
(20100*10) (20100*7)
Contribution margin 10 50% 201000 13 65% 261300
Less: Fixed expenses 146000 146000
Net operating income 55000 115300
c. Yes.Company should automate it's operations since it results in increased net operating income of $ 115300